A 2000L tank containing 550L of water is being filled with water at the rate of 75L/minute from a full 1600L tank. How long will it be before the two tanks have the same amount of water?
Answer by robertb(5830)

Let m = # minutes the 2 tanks will have the same volume. Then
550 + 75m = 1600 - 75m
-1050 = -150m
m = 7.